OpenStudy (anonymous):

these marbles are placed in a bag and two of them are randomly drawn what is the probability of drawing two pink marbles if the first one is placed back in the bag before the second draw thebag has 10 marbles.2 yellow.3 pink.5 blue

OpenStudy (misty1212):

HI!!

OpenStudy (anonymous):

Hi

OpenStudy (misty1212):

the probability the first one is pink is \(\frac{3}{10}\) since there are 3 pink ones and ten total

OpenStudy (misty1212):

since you put it back the probability the second one is pink is also \(\frac{3}{10}\)

OpenStudy (anonymous):

Hi

OpenStudy (misty1212):

multiply those to numbers to get your answer i.e \[\frac{3}{10}\times \frac{3}{10}\]

OpenStudy (anonymous):

9/100?

OpenStudy (anonymous):

Hello?